# Annealed and quenched representations of the Gauss-Rényi measure by “periodic points”

### v1

@@ Abstract.
We consider independently identically distributed random compositions of the Gauss and Rényi maps that generate random continued fractions. Using methods of ergodic theory, thermodynamic formalism and large deviations, we show that weighted cycles of this random dynamical system equidistribute with respect to the Gauss-Rényi measure. We present both annealed (sample-averaged) and quenched (samplewise) results.
@@
Keywords: random dynamical system; periodic points; the Gauss map; the Rényi map
@@ 1. Introduction
It is the essential difference between statistical properties of the sequences $(A_{n}(x))_{n=1}^{\infty}$ and $(B_{n}(x))_{n=1}^{\infty}$ that makes the random continued fraction interesting. For Lebesgue almost every irrational $x$ in $(0,1)$, each positive integer $k$ appears in $(A_{n}(x))_{n=1}^{\infty}$ with frequency $\frac{1}{\log 2}\log\frac{(k+1)^{2}}{k(k+2)}$, while the frequency of $2$ in $(B_{n}(x))_{n=1}^{\infty}$ is $1$. This is due to the fact that $T_{0}$ leaves invariant the Gauss measure $d\lambda_{0}=\frac{1}{\log 2}\frac{dx}{x+1}$, while $T_{1}$ leaves invariant the infinite measure $\frac{dx}{x}$. More precisely, $x=0$ is a neutral fixed point of $T_{1}$: $T_{1}0=0$ and $T_{1}^{\prime}0=1$. For more comparisons of the regular and backward continued fractions as well as more information on the singular behavior of the digit sequence in the backward continued fraction, see [1, 2, 19, 38, 42] for example.
@@ 1.1. Statements of results
We consider an independently identically distributed (i.i.d.) random dynamical system generated by $T_{0}$ and $T_{1}$. This means that $T_{1}$ is chosen with a fixed probability $p\in(0,1)$ at each step. Let $m_{p}$ denote the Bernoulli measure on the sample space $\Omega$ associated with the probability vector (1p,p)(1-p,p). By [18, Theorem 5.2], there exists a unique Borel probability measure $\lambda_{p}$ on $[0,1]$ that is absolutely continuous with respect to the Lebesgue measure on $[0,1]$ and satisfies $\mu=(1-p)\cdot\mu\circ T_{0}^{-1}+p\cdot\mu\circ T_{1}^{-1}$. The measure $\lambda_{p}$, called the Gauss-Rényi measure, is significant since for $m_{p}$-almost every $\omega\in\Omega$ and Lebesgue almost every $x\in\Lambda_{\omega}$, we have
@@
For $p\in[0,1)$, let $h_{p}\colon[0,1]\to[0,\infty)$ denote the Radon-Nikodým derivative of $\lambda_{p}$ with respect to the Lebesgue measure on $[0,1]$. We know that $h_{0}(x)=\frac{1}{\log 2}\frac{1}{x+1}$. For any $p\in(0,1)$, $h_{p}$ is bounded from above and away from 0 [23, Proposition 3.4]. An explicit formula for $h_{p}$ is desired, since it is related to the frequency of digits in the random continued fraction expansion (2.1). Up to present, no algebraic formula for $h_{p}$ is known except for the case $p=0$. Kalle et al. proved that $h_{p}$ is $C^{\infty}$ for any $p\in(0,1)$ [24]. Bousoun et al. [6] obtained a functional-analytic formula for $h_{p}$ for $p\in(0,1)$ sufficiently near 0.
@@ 1.2. Method of proofs of the main results
Owing to the existence of the neutral fixed point of the Rényi map $T_{1}$, for the potential function associated with this countable full shift there exists no Gibbs state. To resolve this difficulty, we construct an appropriate induced system that is topologically conjugate to another countable full shift, and then apply the result of the second-named author in [42]. This requires verifying the regularity of the associated induced potential.
@@ 2.1. A continued fraction algorithm by the Gauss-Rényi map
converges to a number $y\in[0,1]$. For (a) and (b) it suffices to show $x=y$.
@@ 2.2. Large Deviation Principle
Let $X$ be a topological space and let $C(X)$ denote the Banach space of real-valued bounded continuous functions on $X$ endowed with the supremum norm. Recall that the weak* topology on $\mathcal{M}(X)$ is the coarsest topology that makes the map $\mu\in\mathcal{M}(X)\mapsto\int fd\mu$ continuous for any $f\in C(X)$. In this topology, a sequence $(\mu_{n})_{n=1}^{\infty}$ of elements of $\mathcal{M}(X)$ converges to $\mu\in\mathcal{M}(X)$ if and only if $\lim_{n}\int fd\mu_{n}=\int fd\mu$ holds for any $f\in C(X)$. This condition is equivalent to $\lim_{n}\int fd\mu_{n}=\int fd\mu$ for any $f\in C(X)$ that is uniformly continuous (see [36, Chapter 9]).
Donsker and Varadhan have identified three levels of the LDP, see e.g., [12, Chapter I]. The LDP for a sequence of Borel probability measures on $\mathcal{M}(X)$ is referred to as level-2. The LDP for a sequence of Borel probability measures on $\mathbb{R}$ determined by a real-valued function on $X$ is referred to as level-1. By the Contraction Principle, any level-2 LDP can be transferred to a level-1 LDP.
@@ 4.1. Thermodynamic formalism for the countable full shift
Let $\phi\colon\mathbb{N}^{\mathbb{N}}\to\mathbb{R}$ be acceptable and satisfy $P(\phi)<\infty$. Then $\sup\phi$ is finite (see [27, Proposition 2.1.9]). Let
@@
By [27, Theorem 2.1.7], for any $\mu\in\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ we have $h(\mu)+\int\phi d\mu\leq P(\phi)<\infty$, and so $h(\mu)<\infty$. The following equality is known as the variational principle.
Proposition 4.1 ([27, Theorem 2.1.7, Theorem 2.1.8]).
@@
Proposition 4.2 ([27, Theorem 2.2.9, Corollary 2.7.5]).
@@ 4.3. Level-2 LDP for the countable full shift
A proof of this proposition is briefly outline as follows. By the definition (4.5), if $\mu$ is a minimizer of $I_{\phi}$ then there is a sequence $(\mu_{k})_{k=1}^{\infty}$ in $\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ that converges to $\mu$ in the weak* topology with $\lim_{k}F_{\phi}(\mu_{k})=0$. Based on this information we show that $\mu$ is an equilibrium state for $\phi$. The case $\lim_{k}h(\mu_{k})=0$ is easy to handle, while the case $\lim_{k}h(\mu_{k})=\infty$ (and hence $\lim_{k}\int\phi d\mu_{k}\to$-$\infty$) requires attention. A key ingredient in the latter case is the upper semicontinuity of the map $\mu_{k}\mapsto h(\mu_{k})/(-\int\phi d\mu_{k})$, as proved in [40, Theorem 2.4] inspired by [14, Lemma 6.5].
@@
The following proof is almost a repetition of the proof of [40, Theorem 2.1] for the reader’s convenience. Considering $\phi-P(\phi)$ instead of $\phi$, we may assume $P(\phi)=0$. Let $\mu\in\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$ be a minimizer of $I_{\phi}$. Since $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$ is a closed subset of $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$, $\mu$ is shift-invariant. By the definition (4.5), there is a sequence $(\mu_{k})_{k=1}^{\infty}$ in $\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ that converges to $\mu$ in the weak* topology with $\lim_{k}F_{\phi}(\mu_{k})=0$. By [40, Lemma 2.3], we have $\inf_{k}\int\phi d\mu_{k}>$-$\infty$. By this and $\sup\phi<\infty$, a simple upper semicontinuity argument as in [40, Remark 2.5] shows $\int\phi d\mu>$-$\infty$. If $\liminf_{k}h(\mu_{k})=0$, then for any subsequence $(\mu_{k_{j}})_{j=1}^{\infty}$ with $\lim_{j}h(\mu_{k_{j}})=0$ we have
@@
We have $\int\phi d\mu\geq h(\mu)$. If $\int\phi d\mu=0$, then clearly $\mu$ is an equilibrium state for $\phi$. If $\int\phi d\mu>0$, then by [40, Theorem 2.4] we have
@@ 4.4. Symbolic coding of the Gauss-Rényi map
and call $\eta_{p}$ the annealed Gauss-Rényi measure. From (b) and (c) in Proposition 2.1, we have $\Lambda_{\omega}=(0,1)\setminus\mathbb{Q}$ for every $\omega\in\Omega_{0}$. This implies $\Omega_{0}\times((0,1)\setminus\mathbb{Q})\subset\Lambda$, and so $(m_{p}\otimes\lambda_{p})(\Lambda)=1$. Hence $\eta_{p}$ is a probability. The measure $m_{p}\otimes\lambda_{p}$ is $R$-invariant [23, Theorem 3.2] and by [23, Theorem 3.3] it is mixing. Hence $\eta_{p}$ is $\sigma$-invariant and mixing.
@@ 4.5. Refined distortion estimates
The next general lemma on distortions for iterations of an interval map with a neutral fixed point was shown in the proof of [20, Lemma 5.3].
Lemma 4.8 (cf. [20, Lemma 5.3]).
@@ 4.6. Variational characterization of the annealed Gauss-Rényi measure
A proof of Proposition 4.14 breaks into two steps. We first show that $\eta_{p}$ is an equilibrium state for the potential $\psi$. We then establish the uniqueness of equilibrium state for the potential $\psi$. To overcome the lack of regularity of $\psi$ in the second step, we take an inducing procedure that is now familiar in the construction of equilibrium states (see e.g., [27, Section 8], [30]).
@@
Step 2: establishing the uniquness of equilibrium state. Recall that $\widehat{\sigma}\colon{\widehat{\mathbb{N}}}^{\mathbb{N}}\to{\widehat{\mathbb{N}}}^{\mathbb{N}}$ is the induced system associated with the inducing scheme $(\mathbb{N}^{\mathbb{N}}\setminus[1],t_{\mathbb{N}^{\mathbb{N}}\setminus[1]})$ of the left shift $\sigma\colon\mathbb{N}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$ (see §4.2). For the induced potential $\widehat{\psi}$ in (4.18), define $\Psi\colon\mathbb{M}^{\mathbb{N}}\to\mathbb{R}$ by
@@
belongs to $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma),$ and by Abramov-Kac’s formula [30, Theorem 2.3]
@@ Appendix A Periodic continued fractions
By [22, Theorem 1.1] we obtain $x=\lim_{n}r_{n}(x)$, which is the minus continued fraction expansion of $x$:

### v2

@@ Abstract.
We consider independently identically distributed random compositions of the Gauss and Rényi maps that are related to Diophantine approximations. Elaborating on methods in ergodic theory, thermodynamic formalism and large deviations, we show that weighted cycles of this random dynamical system equidistribute with respect to the Gauss-Rényi measure. We present both annealed (sample-averaged) and quenched (samplewise) results.
@@
Keywords: random dynamical system; periodic points; the Gauss map; the Rényi map; thermodynamic formalism; large deviations
@@ 1. Introduction
It is the essential difference between statistical properties of the sequences $(A_{n}(x))_{n=1}^{\infty}$ and $(B_{n}(x))_{n=1}^{\infty}$ that makes the random continued fraction interesting. For Lebesgue almost every irrational $x$ in $(0,1)$, each positive integer $k$ appears in $(A_{n}(x))_{n=1}^{\infty}$ with frequency $\frac{1}{\log 2}\log\frac{(k+1)^{2}}{k(k+2)}$, while the frequency of $2$ in $(B_{n}(x))_{n=1}^{\infty}$ is $1$. This is due to the fact that $T_{0}$ leaves invariant the Gauss measure $d\lambda_{0}=\frac{1}{\log 2}\frac{dx}{x+1}$, while $T_{1}$ leaves invariant the infinite measure $\frac{dx}{x}$. More precisely, $x=0$ is a neutral fixed point of $T_{1}$: $T_{1}0=0$ and $T_{1}^{\prime}0=1$. For more comparisons of the regular and backward continued fractions as well as more information on the singular behavior of the digit sequence in the backward continued fraction, see [1, 2, 20, 21, 35, 41, 45] for example.
@@ 1.1. Statements of results
We consider an independently identically distributed (i.i.d.) random dynamical system generated by $T_{0}$ and $T_{1}$. This means that $T_{1}$ is chosen with a fixed probability $p\in(0,1)$ at each step. Let $m_{p}$ denote the Bernoulli measure on the sample space $\Omega$ associated with the probability vector (1p,p)(1-p,p). By [19, Theorem 5.2], there exists a unique Borel probability measure $\lambda_{p}$ on $[0,1]$ that is absolutely continuous with respect to the Lebesgue measure on $[0,1]$ and satisfies $\mu=(1-p)\cdot\mu\circ T_{0}^{-1}+p\cdot\mu\circ T_{1}^{-1}$. The measure $\lambda_{p}$, called the Gauss-Rényi measure, is significant since for $m_{p}$-almost every $\omega\in\Omega$ and Lebesgue almost every $x\in\Lambda_{\omega}$, we have
@@
For $p\in[0,1)$, let $h_{p}\colon[0,1]\to[0,\infty)$ denote the Radon-Nikodým derivative of $\lambda_{p}$ with respect to the Lebesgue measure on $[0,1]$. We know that $h_{0}(x)=\frac{1}{\log 2}\frac{1}{x+1}$. For any $p\in(0,1)$, $h_{p}$ is bounded from above and away from 0 [25, Proposition 3.4]. An explicit formula for $h_{p}$ is desired, since it is related to the frequency of digits in the random continued fraction expansion (2.1). Up to present, no algebraic formula for $h_{p}$ is known except for the case $p=0$. Kalle et al. proved that $h_{p}$ is $C^{\infty}$ for any $p\in(0,1)$ [26]. Bahsoun et al. [6] obtained a functional-analytic formula for $h_{p}$ for $p\in(0,1)$ sufficiently near 0.
@@ 1.2. Method of proofs of the main results
Owing to the existence of the neutral fixed point of the Rényi map $T_{1}$, for the potential function associated with this countable full shift there exists no Gibbs state. To resolve this difficulty, we construct an appropriate induced system that is topologically conjugate to another countable full shift, and then apply the result of the second-named author in [45]. This requires verifying the regularity of the associated induced potential. The use of induced systems for an analysis of random dynamical systems with infinitely many branches can be found, for example, in [11].
@@ 2.1. A continued fraction algorithm by the Gauss-Rényi map
converges to a number $y\in[0,1]$. Moreover, $y\notin\mathbb{Q}$ if and only if $(-1)^{\omega_{n+1}}+C_{n}(\omega,x)\geq 2$ for infinitely many $n\in\mathbb{N}$. Hence, for (a) and (b) it suffices to show $x=y$.
@@ 2.2. Large Deviation Principle
Let $X$ be a topological space and let $C(X)$ denote the Banach space of real-valued bounded continuous functions on $X$ endowed with the supremum norm. Recall that the weak* topology on $\mathcal{M}(X)$ is the coarsest topology that makes the map $\mu\in\mathcal{M}(X)\mapsto\int fd\mu$ continuous for any $f\in C(X)$. In this topology, a sequence $(\mu_{n})_{n=1}^{\infty}$ of elements of $\mathcal{M}(X)$ converges to $\mu\in\mathcal{M}(X)$ if and only if $\lim_{n}\int fd\mu_{n}=\int fd\mu$ holds for any $f\in C(X)$. This condition is equivalent to $\lim_{n}\int fd\mu_{n}=\int fd\mu$ for any $f\in C(X)$ that is uniformly continuous (see [39, Chapter 9]).
Donsker and Varadhan have identified three levels of the LDP, see e.g., [13, Chapter I]. The LDP for a sequence of Borel probability measures on $\mathcal{M}(X)$ is referred to as level-2. The LDP for a sequence of Borel probability measures on $\mathbb{R}$ determined by a real-valued function on $X$ is referred to as level-1. By the Contraction Principle, any level-2 LDP can be transferred to a level-1 LDP.
@@ 4.1. Thermodynamic formalism for the countable full shift
Let $\phi\colon\mathbb{N}^{\mathbb{N}}\to\mathbb{R}$ be acceptable and satisfy $P(\phi)<\infty$. Then $\sup\phi$ is finite (see [29, Proposition 2.1.9]). Let
@@
By [29, Theorem 2.1.7], for any $\mu\in\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ we have $h(\mu)+\int\phi d\mu\leq P(\phi)<\infty$, and so $h(\mu)<\infty$. The following equality is known as the variational principle.
Proposition 4.1 ([29, Theorem 2.1.7, Theorem 2.1.8]).
@@
Proposition 4.2 ([29, Theorem 2.2.9, Corollary 2.7.5]).
@@ 4.3. Level-2 LDP for the countable full shift
A proof of this proposition is briefly outline as follows. By the definition (4.5), if $\mu$ is a minimizer of $I_{\phi}$ then there is a sequence $(\mu_{k})_{k=1}^{\infty}$ in $\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ that converges to $\mu$ in the weak* topology with $\lim_{k}F_{\phi}(\mu_{k})=0$. Based on this information we show that $\mu$ is an equilibrium state for $\phi$. The case $\lim_{k}h(\mu_{k})=0$ is easy to handle, while the case $\lim_{k}h(\mu_{k})=\infty$ (and hence $\lim_{k}\int\phi d\mu_{k}\to$-$\infty$) requires attention. A key ingredient in the latter case is the upper semicontinuity of the map $\mu_{k}\mapsto h(\mu_{k})/(-\int\phi d\mu_{k})$, as proved in [43, Theorem 2.4] inspired by [15, Lemma 6.5].
@@
The following proof is almost a repetition of the proof of [43, Theorem 2.1] for the reader’s convenience. Considering $\phi-P(\phi)$ instead of $\phi$, we may assume $P(\phi)=0$. Let $\mu\in\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$ be a minimizer of $I_{\phi}$. Since $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$ is a closed subset of $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma)$, $\mu$ is shift-invariant. By the definition (4.5), there is a sequence $(\mu_{k})_{k=1}^{\infty}$ in $\mathcal{M}_{\phi}(\mathbb{N}^{\mathbb{N}},\sigma)$ that converges to $\mu$ in the weak* topology with $\lim_{k}F_{\phi}(\mu_{k})=0$. By [43, Lemma 2.3], we have $\inf_{k}\int\phi d\mu_{k}>$-$\infty$. By this and $\sup\phi<\infty$, a simple upper semicontinuity argument as in [43, Remark 2.5] shows $\int\phi d\mu>$-$\infty$. If $\liminf_{k}h(\mu_{k})=0$, then for any subsequence $(\mu_{k_{j}})_{j=1}^{\infty}$ with $\lim_{j}h(\mu_{k_{j}})=0$ we have
@@
We have $\int\phi d\mu\geq h(\mu)$. If $\int\phi d\mu=0$, then clearly $\mu$ is an equilibrium state for $\phi$. If $\int\phi d\mu>0$, then by [43, Theorem 2.4] we have
@@ 4.4. Symbolic coding of the Gauss-Rényi map
and call $\eta_{p}$ the annealed Gauss-Rényi measure. From (b) and (c) in Proposition 2.1, we have $\Lambda_{\omega}=(0,1)\setminus\mathbb{Q}$ for every $\omega\in\Omega_{0}$. This implies $\Omega_{0}\times((0,1)\setminus\mathbb{Q})\subset\Lambda$, and so $(m_{p}\otimes\lambda_{p})(\Lambda)=1$. Hence $\eta_{p}$ is a probability. The measure $m_{p}\otimes\lambda_{p}$ is $R$-invariant [25, Theorem 3.2] and by [25, Theorem 3.3] it is mixing. Hence $\eta_{p}$ is $\sigma$-invariant and mixing.
@@ 4.5. Refined distortion estimates
The next general lemma on distortions for iterations of an interval map with a neutral fixed point was shown in the proof of [22, Lemma 5.3].
Lemma 4.8 (cf. [22, Lemma 5.3]).
@@ 4.6. Variational characterization of the annealed Gauss-Rényi measure
A proof of Proposition 4.14 breaks into two steps. We first show that $\eta_{p}$ is an equilibrium state for the potential $\psi$. We then establish the uniqueness of equilibrium state for the potential $\psi$. To overcome the lack of regularity of $\psi$ in the second step, we take an inducing procedure that is now familiar in the construction of equilibrium states (see e.g., [29, Section 8], [32]).
@@
Step 2: establishing the uniqueness of equilibrium state. Recall that $\widehat{\sigma}\colon{\widehat{\mathbb{N}}}^{\mathbb{N}}\to{\widehat{\mathbb{N}}}^{\mathbb{N}}$ is the induced system associated with the inducing scheme $(\mathbb{N}^{\mathbb{N}}\setminus[1],t_{\mathbb{N}^{\mathbb{N}}\setminus[1]})$ of the left shift $\sigma\colon\mathbb{N}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$ (see §4.2). For the induced potential $\widehat{\psi}$ in (4.18), define $\Psi\colon\mathbb{M}^{\mathbb{N}}\to\mathbb{R}$ by
@@
belongs to $\mathcal{M}(\mathbb{N}^{\mathbb{N}},\sigma),$ and by Abramov-Kac’s formula [32, Theorem 2.3]
@@ Appendix A Periodic continued fractions
By [24, Theorem 1.1] we obtain $x=\lim_{n}r_{n}(x)$, which is the minus continued fraction expansion of $x$:
@@ Acknowledgments
Data Availability
This article has no associated data and material.
Conflict of interest
The authors have no conflicts of interest to declare that are relevant to the content of this article.